OpenStudy (anonymous):

integrate(0 to 1) x cos ((2n-1)pi x/2)sqrt 2 dx

OpenStudy (anonymous):

\[\int\limits_{0}^{1}x \cos(\frac{ (2n-1)\pi x }{ 2 })\sqrt{2}dx\]